TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-07-12 23:31:00
subject: 7\08 Pt 1 ISS On-Orbit Status 08-07-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

9 July 2003

ISS On-Orbit Status 8 July 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.  Day 75 in space for the crew.

Using the MO-21 "Ecosfera" air sampler and incubation equipment,
broken out and set up yesterday, CDR Malenchenko spent about 40
minutes collecting air samples for atmospheric microbial analysis.
[MO-21 determines microbial contamination of the ISS atmosphere,
specifically the total bacterial and fungal microflora counts and
microflora composition according to morphologic criteria of
microorganism colonies.] 

Malenchenko then worked on restoring the Russian laptop 1 to
functionality by first creating a boot-up floppy disk (for Norton
Ghost 5.1D) on Russian laptop 2.  Later, he activated laptop 1,
installed a software patch and "cloned" a new HDD (hard disk drive)
on the machine from the Norton Ghost 5.1D disk obtained from laptop
2. 

FE/SO Ed Lu meanwhile spent time in the U.S. segment (USOS) with the
regular periodic inspection of smoke detectors in the Lab module,
Airlock (A/L) and Node "Unity", as well as of the A/L's microbial
filter. 

Yuri Malenchenko completed the scheduled installation of new
satellite navigation equipment (ASN), delivered on Progress 11P, in
the Service Module (SM).  [The ASN was connected to the onboard
computer system's laptop 3, configured for the subsequent test.
After he had mated BITS2-12 onboard measurement telemetry system
connectors to the ASN, a raw data transfer test to laptop 3 was to be
conducted from TsUP by remote control.] 

The CDR attended to the routine maintenance of SOZh life support
systems, while Ed Lu prepared the daily IMS (inventory management
system) "delta" file for updating the IMS database and completed the
regular daily status checkup of the autonomous Lab payloads..

Ed also attended to the weekly transfer of accumulated data files
with workout data from the TVIS and RED exercise equipment to the MEC
via PCMCIA cards (personal computer memory card international
association) for subsequent downlink to Earth.

At 5:45am EDT, both crewmembers participated in an interactive
televised PAO/educational event sponsored by NASDA, with Japanese
middle school students gathered at the Tokyo FM Radio station banquet
hall (Ku- & S-band). Former NASDA Astronaut Mamoru Mohri attended.

Both crewmembers completed their daily physical exercise program
(2.5 hrs.) on TVIS treadmill, CEVIS cycle ergometer, and RED
anaerobic exerciser. 

The Science Officer changed lenses of the EarthKAM (EK) camera at the
Lab nadir window from 50mm to 180mm.  EK team reported an excellent
startup yesterday, and the on-board system is working well.
[Attending today's and tomorrow's National Geographic
Society-sponsored sessions are 25 students, ranging from 5 th grade
to college sophomores.  Commented one of their teachers:  "I can't
think of another program that interests and excites students to the
point that they are willing to come in during vacation (some home
from college), to work during a mission."  EarthKAM (Earth Knowledge
Acquired by Middle school students) employs an ESC 460C electronic
still camera mounted at the Lab's down-facing window, normally
equipped with a 50mm lens (f1.4).  Powered by 16Vdc from a 120 Vdc
adapter, EarthKAM takes pictures by remote operation from the ground,
without crew interaction.  It is available for Middle School-ers, who
submit image requests and conduct geographic research.  The requests
are uplinked in a camera control file to an SSC (station support
computer) laptop which then activates the camera at specified times
and receives the digital images from the camera's storage card on its
hard drive, for subsequent downlink via OpsLAN.] 

The Kodak 760 ECS (electronic still camera) used by Ed Lu to
photograph stunning aurora images has shown some noise, due to
"pixelation" or "hot pixels" not properly removed or
blended into the
image by the automatic camera procedure.  The most probable cause is
excessive radiation as ISS passed through the aurora.  All noise will
be removed from the photos by ground processing.

Today's CEO targets, no longer limited in the current LVLH attitude
and including the targets of the Lewis & Clark 200-year memorial
locations, were Baghdad, Iraq (nadir pass), Phoenix, AZ (NIGHT
TARGET: nadir pass), Lisbon, Portugal (nadir pass), Pomeroy,
Washington (LEWIS & CLARK SITE:  Nadir pass--looking on the south
side of the Snake River [within the second bend of the river upstream
of its confluence with the Columbia River]), Lemhi Pass, Mont/Idaho
(LEWIS & CLARK SITE:  Nadir pass--shooting mountain range crests
[dark forest green, as opposed to lighter, browner river valleys] and
the crew should have gotten this now-unused pass over the continental
divide.  The explorers crossed here into Idaho on August 12, 1805.
Lemhi Pass was abandoned for the westward migrations since other
lower passes exist), Denver, Colorado (nadir pass), and Los Angeles,
California (nadir and a touch left).

CEO images can be viewed at the websites
http://eol.jsc.nasa.gov
http ://earthobservatory.nasa.gov
See also the website "Space Station Challenge" at
http://voyager.cet.edu/iss/

U.S. and Russian Segment Status (as of 1:30pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS): Elektron O2 generator is powered On (16 amp mode).  Vozdukh
CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is
Off.  TCCS (trace contaminant control subsystem) is operating.  MCA
(major constituents analyzer) is operating.  BMP Harmful Impurities
unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode (unit is
off).  RS air conditioner SKV-1 is Off; SKV-2 is Off.

SM Working Compartment:   Pressure (mmHg) -- 742; temperature (deg C)
-- 27.5; ppO2 (mmHg) -- data invalid ; ppCO2 (mmHg) -- data invalid .
SM Transfer Compartment:  Pressure (mmHg) -- 751; temperature (deg C)
-- 20.4.
FGB Cabin:   Pressure (mmHg) -- 752; temperature (deg C) -- n/a.
Node: Pressure (mmHg) -- 741.34; temperature (deg C) -- 23.0 (shell);
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab:   Pressure (mmHg) -- 743.40; temperature (deg C) -- n/a;
ppO2 (mmHg) -- 169.1; ppCO2 (mmHg) -- 3.8.
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 743.50; temperature
(deg C) -- n/a; shell heater temp (deg C) -- 22.3, ppO2 (mmHg) --
n/a; ppCO2 (mmHg) -- n/a.
PMA-1:   Shell heater temp (deg C) -- 23.9
PMA-2:  Shell heater temp (deg C) -- 20.2.
         (n/a = data not available)

Propulsion System (PS):
Total propellant load available: 3803 kg (8384 lb) as of 6/26
[SM(774) + FGB(2447) + Progress M(182) +Progress M-1(400)].
(Capability: SM -- 860 kg; FGB -- 6120 kg).

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.